OpenStudy (anonymous):

What is the approximate measure of the diagonal of a rectangle with dimensions 17 cm by 4 cm? a. 4.6 cm b. 16.4 cm c. 16.5 cm d. 17.5 cm

OpenStudy (anonymous):

17.5

Parth (parthkohli):

 This forms a right-triangle. Using the pythagorean theorem: \(\Large \color{MidnightBlue}{\Rightarrow a^2 + b^2 = c^2 }\) c is the diagonal. \(\Large \color{MidnightBlue}{\Rightarrow 17^2 + 4^2 = c^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow 289 + 16 = c^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow 305 = c^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow c \approx 17.5 }\)